OpenStudy (anonymous):

A pyramid has a volume of 24 and a base area of 12. What is the length of the altitude of the pyramid ?

OpenStudy (campbell_st):

The volume of a pyramid is \[V = \frac{1}{3} \times ..base.area \times height\] you know the base area and the volume so \[24 = \frac{1}{3} \times 12 \times height\] just solve for height